OpenStudy (anonymous):

Find the inverse of the function. f(x) = x^3 - 4 ??

OpenStudy (mertsj):

y=x^3-4 x=y^3-4 Solve for y

OpenStudy (agent0smith):

y = x^3 - 4 swap x and y like so x = y^3 - 4 Now rearrange this to get y = ...